If an electron is accelerated from rest through a potential difference of 1500 V, what speed does it reach? (e = 1.60x10^-19 C , mass electron = 9.11x 10^-31 kg)

A. 1.1 x 10^7 m/s

B. 1.9 x 10^7 m/s

C. 1.5 x 10^7 m/s

D. 2.3 x 10^7 m/s
Answer #1

energy gained due to potential difference = 1.6 * 10-19 * 1500 = KE

0.5 * 9.1 * 10-31 * V^2 = 1.6*10-19 * 1500

V = 2.3 * 107 m/s